You are given the following utility function and price of commodities q1 and q2:

U = 3q1+q1q2-5q2-15

P1=3 and p2=2

If the corresponding bugdet is 20.

i. Write the consumer's budget equation

ii.construct a constrained utility maximization problem out of the information given above

iii. Write the augmented objective function?

Iv. Find the optimum level of U

V. Is the second order condition for a maximum satisfied?

Vi. Find the levels of q1 and q2 that will satisfy the first order condition for a maximum.


1
Expert's answer
2021-07-19T16:11:54-0400

Given

U = 3q1+q1q2-5q2-15

P1=3 and p2=2

Utility is maximized at the point where indifference curves are tangential to the budget constraint .  

i) Budget Constraint : 

There are two sides to BC expenditure side and the income side . 

Expenditure side "= P_1q_1 + P_2q_2"

Income Side"= m = 20"

Budget Constraint :

"P_1q_1 + P_2q_2 =< 20"


ii) Utility maximization problem is optimized where : MRS ( marginal rate of substitution )"= \frac{P_1}{P_2}"



iii) Objective function is referred to the function which is to be optimized subject to the constraint .

Here utility function is to be maximized :

"U_{max} = 3q_1+q_1q_2-5q_2-15"

"s.t P_1q_1 + P_2q_2 =< 20"


iv) Optimization point is where :

"MRS = \frac{P_1}{P_2}\\

MRS =\frac{ MU_{q1}}{MU_{q2}} \\

MU_{q1} = u'(q_1) = 3 + q_2\\

MU_{q2} = u'(q_2) = q_1 - 5"


v)

Second Order Condition of the utility maximization problem is that the marginal utility of of any good should be diminishing .

"U''(q_1)" should less than or equal to 0. 

"U''(q_1 ) = \frac{du^2}{(dq_1)^2} = \frac{d(3 + q_2)}{dq_1} = 0"

Hence SOC is justified .


vi

Putting in optimization equation :

"\frac{(3 + q_2)}{(q_1 - 5 )} = \frac{P_1}{ P_2}\\

\frac{(3 + q_2)}{(q_1 - 5 )} = \frac{ 3}{2}\\

6 + 2q_2 = 3q_1 - 15 \\

3q_1 - 2q_2 = 21 --------(i)

also\space budget\space constraint\space is \space given \space as :\\

P_1q_1 + P_2q_2 = 20 \\

3q_1 + 2q_2 = 20 --------(ii)\\

Solving\space equation\space (i) \space and\space (ii)\space simultaneously\space using\space elimination\space we\space get :\\

6q_1 = 41 \\

q_1 = \frac{41}{6 }= 6.83\\

Putting\space this\space into\space (i)\\

q_2 = 2.44"